OpenStudy (anonymous):

The ratio of boys to girls at a school is 12 to 13. If 300 students attend the school, how many are boys?

OpenStudy (anonymous):

boys = 12/(12+13) * 300 = ...

OpenStudy (anonymous):

12x + 13x = 25x not 15x, @Luis_Rivera